OpenStudy (anonymous):

What are the steps to solve this: 3m-9n=24; n= -1,1,3

OpenStudy (anonymous):

Is this the exact question from the book or hw assignment? The written part not the expression itself

OpenStudy (anonymous):

hw assignment, exactly written.

OpenStudy (anonymous):

int this case just plug in the different n values and solve for m there should be 3 different m values

OpenStudy (anonymous):

in*

OpenStudy (anonymous):

you understood how to solve for the problem?

OpenStudy (anonymous):

kind of.

OpenStudy (anonymous):

3m - 9n =24 when n = -1 plug in n into the equation and you get 3m - 9(-1) = 24 => 3m + 9 = 24 now solve for m this means try and get m by itself without any coefficient (number in the front) so: 3m+ 9-9 = 24-9 => subtracting 9 on both sides you now have 3m = 15 at this point you have to get rid of the 3 on the left hand side how to do this is multiplying by the reciprocal which is 1/3 m = 3. That is just for when n = -1 Now do the same steps for when n = 1 & n = 3

OpenStudy (anonymous):

okay I get it now thanks!